OpenStudy (anonymous):

Given the quadratic function f(x) = 9x2 – 36x + 36, find a value of x such that f(x) = 9.

OpenStudy (mertsj):

[9x^2-36x+36=9\] \[9x^2-36x+27=0\]\ \[x^2-4x+3=0\]

OpenStudy (mertsj):

Factor and solve

OpenStudy (maheshmeghwal9):